(a) What is the pKbof NO2?

(

Ka of nitrous acid is 7.1

×

10−4

)



pKb=

(b) What is the pKaof the hydrazinium ion, H2N-NH3+?

(

Kb of hydrazine = 8.5

×

10−7

)



pKa=
Answer #1

a) HNO2 ka = 7.1 x 10-4

pKa = - log Ka = - log [7.1 x 10-4] = 3.15

pKa + pKb = 14

pKb = 14 - 3.15

pKb = 10.85

b) NH2-NH2 = Kb = 8.5 x10-7

pKb = - log Kb = - log [8.5 x 10-7] = 6.07

pKb + pKa = 14

pKa = 14 - 6.07

pKa = 7.93